|
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
|
|
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
|
|
SASE1 Status and SASE3 Readiness
Technical integration and coordination
|
| Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
|
Message ID: 579
Entry time: 19 Nov 2018, 10:39
|
| Author: |
Adriano Violante |
| Location: |
General |
| Severity: |
Information |
| Subject: |
Status report after dispatch meeting 09.30 |
|
|
SPB
- Gotthard needs to be tested.
- Pulse picker will be used and commissioning still has to be done.
- Sample chamber encoder drift issue. Has to be investigated further.
FXE
SQS
- Beam during last WE: no major issues
- Interlocks will be updated tomorrow
- Some issues concerning direction of motors need to be fixed with AE.
- AVC300 valve. Firmware will be available this week.
- Digitizers and photonics camera are needed. Work ongoing
SCS
- Motors for the Fast CCD cameras are not working properly and interlock about pressure for the detector is needed ==> Loop 7 needs to be updated, AE will take care of that.
- When the CCD will inject data into the DAQ ITDM will be informed.
- There was a request to CAS to provide values for the Monochromator at a rate faster that 10 Hz. This requires evaluation of the feasibility by the experts (CAS/AE). For the time being the middle layer will not be changed.
MID
- AGIPD commissioning:
- Loop 11 needed. Today tests by AE. Until Tue. evening all the intelock tests should be done. From Wed. on operation.
- uTCA crate needs to be operational too. According to AE can be done.
- Thu. on: DQA tests.
HED
XRO
- Temporary solution concerning Karabo topics management to alllow the reading of the Monochromator values by SCS. The definitive fix will come in the future.
XPD
- XTD10 XGM repaired
- XTD1 ph. science camera was replaced with the one previously installed in SASE1. Repair ongoing.
- the company went bankrupt and was bought by a new company. For info contact A. Koch.
VAC
AE
- OCD: intervention request during the WE to fix a train id issue. Should be ok now but the problem still needs to be reproduced and undertstood.
ITDM
CAS
- Release tests with ITDM ongoing.
AoB |